# Supplemental Material to ”The Emergence of Vertical Diversity under Disturbance”

### v2

@@ I The equations to determine the stationary solutions of the population dynamics
$V_{h}({\bm{Q}})=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
@@
Let us remind that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have
@@ I.1 The explicit expressions of $Q_{h}^{*}$ and $\chi_{h}$ at $r\to 0$
$Q_{2}^{*}(T,0)=\frac{\exp(-2\beta)}{2+\exp(-2\beta)}.$
@@
There are $6$ cases in terms of the possibilities of taking values of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$, each of which can be $-1$, 00, or $1$. Note that if one of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$ is 00, the value of the other is uniquely determined. Indeed, when the two conditions of
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1>0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (9)
@@
for $h=1,3$. Additionally, the two conditions of (9) lead to
@@
$Q_{3}^{*}(0,0)=1-Q_{1}^{*}(0,0),$
$Q_{2}^{*}(0,0)=0.$
Note that the other $5$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
@@
Then, substituting this form with (7) into (4), for $O(\beta r)\ll 1$ up to $O(r)$ order, we obtain
@@
$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$
@@
I.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
Among $6$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, let us consider the situation where the following two conditions hold for $T\ll 1$:
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1=0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0.\end{cases}$ (18)
@@ I.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
Combining (21) with the first condition of (18) gives
@@
$Z_{3}=$ $Q_{3}^{*}(0,r)=-\frac{1}{4}r+\frac{1}{4},$
$Q_{2}^{*}(0,r)=\frac{r}{2}.$
Let us move onto (20) into which we substitute (23), leading to
@@
Since $-1<\tanh 2(-\mu_{1}+\mu_{3})<1$ should hold, $r$ satisfies $0<r<r_{0}=-3+2\sqrt{3}$ due to $\lim_{r\to 0}\theta(r)=1$ and $\lim_{r\to r_{0}}\theta(r)=-1$. Thus, it turns out that (23) as the stationary solution is valid only for $0<r<r_{0}$. Note that $\lim_{r\to 0}Q_{1}^{*}(0,r)=\sup Q_{1}^{*}(0,0)=\frac{3}{4}$, $\lim_{r\to 0}Q_{3}^{*}(0,r)=\inf Q_{3}^{*}(0,0)=\frac{1}{4}$, and $\lim_{r\to 0}Q_{2}^{*}(0,r)=Q_{2}^{*}(0,0)=0$ hold.
Next, among the other $5$ possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, when the following two conditions
$\begin{cases}-2Q_{1}^{*}+2Q_{3}^{*}+1<0,\\ 2Q_{1}^{*}-2Q_{3}^{*}+1>0,\end{cases}$ (25)
@@
$Q_{3}^{*}(0,r)=\frac{(1-r)^{2}}{(1+r)^{2}},$
$Q_{2}^{*}(0,r)=\frac{2r(1-r)}{(1+r)^{2}}.$
Substituting (29) into the two conditions of (25) gives
$\left\{\begin{array}[]{l}3r^{2}+10r-1>0,\\ r^{2}+6r-3>0,\end{array}\right.$
which lead to $r_{0}<r\leq 1$. Thus, the obtained solution (29) is valid only for $r_{0}<r\leq 1$. Combining two cases of smaller $r$ and larger $r$ than $r_{0}$, the left-sided limit and the right-sided limit of $Q_{h}^{*}$ as $r\to r_{0}$ coincide. Note that the other $4$ cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of equation (4).
@@
$\mu_{3}(0,r)=\frac{1-r}{4}\text{artanh }\theta(r).$
@@
I.3 The explicit expressions of $Q_{h}^{*}$ and $\nu_{h}$ at $\beta\to 0$
@@
$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$
$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3},$
which are shown in Fig. 1(a).
(a)Refer to caption
(b)Refer to caption
Figure 1: (a) $Q_{h}^{*}$ as a function of $r$ for $T=\infty$. (b) $D_{\alpha}$ as a function of $r$ for $T=\infty$.
@@
Note that by taking the derivative of the stationary solutions from (39) with respect to $r$, we also have
@@
$\chi_{2}(\infty,0)=-\frac{-4}{9},$ (45)
which are consistent with the expression of (17) in the limit of $\beta\to 0$.
II Hill numbers $D_{\alpha}(r)$ as a function of $r$
Let us define $M_{\alpha}:=\log D_{\alpha}$ and ${\bm{Q}}^{*}:=(Q_{1}^{*},Q_{2}^{*},Q_{3}^{*})\in[0,1]^{3}$. Then, the sign of $\partial_{r}M_{\alpha}$ is the same as that of $\partial_{r}D_{\alpha}$. Indeed, we get the following:
$F_{\alpha}({\bm{Q}}^{*})=\sum_{h}(Q_{h}^{*})^{\alpha-1}\chi_{h}.$ (47)
II.1 The existence of a local maximum of $D_{\alpha}$ at $T\to 0$ for any $0<\alpha<\infty$
First, let us consider the case of $r<r_{0}$ and $T\to 0$ where we use equations (23) by putting $r=r_{0}$. Then, as shown in Fig. 2(a), we obtain that $F_{\alpha}<0$ for $\alpha>1$, $F_{\alpha}>0$ for $10.199\leq\alpha<$ 1, and $F_{\alpha}<0$ for $\alpha\leq 0.198$ as $r\to r_{0}$. It means that for $\alpha\geq 0.199$, $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is positive, whereas it is negative for $0<\alpha\leq 0.198$.
Second, let us consider the case of $r>r_{0}$ where we use equations (29) by putting $r=r_{0}$. Then, as shown in Fig. 2(b), it turns out that $F_{\alpha}>0$ for $\alpha>1$ and $F_{\alpha}<0$ for $\alpha<1$. It means that $\partial_{r}D_{\alpha}$ as $r\to r_{0}$ is negative for any $\alpha>0$.
(a)Refer to caption
(b)Refer to caption
Figure 2: (a) $\lim_{r\to r_{0}}F_{\alpha}$ as a function of $\alpha$ for $r<r_{0}$. (b) $\lim_{r\to r_{0}}F_{\alpha}$ a function of $\alpha$ for $r>r_{0}$
II.2 The absence of local maxima of $D_{\alpha}$ at $T\to\infty$ for $0<\alpha<\infty$
Suppose the following condition that for a fixed $T$, $\chi_{1}>0$, $\chi_{2}<0$, $\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold for $0<r\leq 1$. On that condition, we can derive that $(Q_{1}^{*})^{\alpha-1}\chi_{1}$ as a term in $F_{\alpha}$ can be larger than $|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ for $\alpha>1$, and smaller than it for $10<\alpha<$ 1. This leads to that for $0<r\leq 1$, $F_{\alpha}>0$ holds for $\alpha>1$, and $F_{\alpha}<0$ holds for $10<\alpha<$ 1, corresponding to that there is no local maximum of $D_{\alpha}$ as a function of $r$ in $0<r\leq 1$ for $\alpha>0$.
$|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|$ $=(Q_{2}^{*})^{\alpha-1}\chi_{1}+(Q_{3}^{*})^{\alpha-1}|\chi_{3}|-(Q_{2}^{*})^{\alpha-1}|\chi_{3}|$
where we have used $\sum_{h}\chi_{h}=0$. Hence, we obtain
$\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}\leq|(Q_{2}^{*})^{\alpha-1}\chi_{2}+(Q_{3}^{*})^{\alpha-1}\chi_{3}|\leq\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1}.$ (49)
It means that for $\alpha>1$, the following holds:
$\chi_{1}(Q_{1}^{*})^{\alpha-1}>\chi_{1}\max(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (50)
$\chi_{1}(Q_{1}^{*})^{\alpha-1}<\chi_{1}\min(Q_{2}^{*},Q_{3}^{*})^{\alpha-1},$ (51)
For $T=\infty$, (39) directly means that the conditions of $\chi_{1}>0,\chi_{2}<0,\chi_{3}<0$, and $Q_{1}^{*}>Q_{2}^{*},Q_{3}^{*}$ hold. Thus, it leads to the absence of local maxima of $D_{\alpha}$ in $0<r\leq 1$ for any $\alpha>0$ as shown in Fig. 1(b).
Note that the conditions mentioned above can be broken, in particular, for sufficiently small $T$, for which the above results are not applied. Nevertheless, using a nullcline method to compute the stationary solutions of (4), we have numerically estimated that there is $\alpha_{0}(T)>0$ such that for $\alpha>\alpha_{0}(T)$, there are no local maxima of $D_{\alpha}$ for $T\geq 0.26$, whereas for $T\leq 0.25$, there is a local maximum of $D_{\alpha}$ for any $\alpha>0$.
III Monte Carlo sampling
III.1 Updating rules
@@
  2. 1.
    Choose randomly a name $\ell\ (1\leq\ell\leq N)$ of an agent with uniform distribution. That is, the probability of taking each name is $1/N$.
  3. 2.
@@
  4. 3.
    If the chosen agent $\ell$ is not disturbed, choose either positive or negative direction with probability $1/2$.
@@
      If positive direction is chosen and $x_{\ell}\neq 3$, make the transition from state $x_{\ell}$ to $x_{\ell}+1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh{\beta\Big{(}2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})}+1\Big{)}\Bigg{)}.$ (52)
@@
      If negative direction is chosen and $x_{\ell}\neq 1$, make the transition from state $x_{\ell}$ to $x_{\ell}-1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh\beta\Big{(}-2(x_{\ell}-\frac{1}{N}\sum_{k\neq\ell}x_{k})+1\Big{)}\Bigg{)}.$ (53)
As shown in Fig. 3, we have confirmed that Monte Carlo simulations have good agreement with the solution of the population dynamics (1).
Figure 3: The Monte Carlo (MC) simulation with $N=1000$ and the solutions of the population dynamics (1). $T=1$, $r=0.2$.
III.2 Finite-size fluctuations
@@
Figure 4: The variance $\sigma_{hh}$ and correlation coefficients $C_{hh^{\prime}}$ averaged over $10$ samples with different trajectories. $T=10^{-3}$, $t_{0}=100$, $\tau=150$, $N=10^{2},10^{3},10^{4}$.
As shown in Fig. 4, we have performed Monte Carlo simulations with different systems size $N$. If we straightforwardly extrapolate the behaviors for larger system size, there seem to be singular behaviors in the three types of variance and the three types of correlation function in the large size limit near $r=r_{0}$ for $r<r_{0}$. In particular, it is notable to point out that three correlation coefficients in the large size limit seem to have perfect correlation, corresponding to $|C_{hh^{\prime}}|=1$, for only $r<r_{0}$.
Thus, the results obtained by Monte Carlo simulations suggest that there is a universal curve presumably independent of $N$. Further, in the larger system sizes, it is more obvious that $C_{hh^{\prime}}$ for $r<r_{0}$ show U-shaped curves. It is notable that the bottoms of the U-shaped curves are close to $r=-5+2\sqrt{7}\simeq 0.291\cdots$, which is the point where $\mu_{h}$ changes its sign. Nevertheless, more works with sufficiently large samples need to be done to obtain more convincing data, which remain to be found as another work in the future.

### v4

# Emergence of vertical diversity under disturbance

@@ Appendix A The equations to determine the stationary solutions of the population dynamics
$V_{h}({\bm{Q}})=-\delta_{h,3}rQ_{3}+Q_{2}U\Bigl{(}2\to h\mid{\bm{Q}}\Bigr{)}$
$+\delta_{h,1}r\sum_{h^{\prime}=2}^{3}Q_{h^{\prime}}-Q_{h}U\Bigl{(}h\to 2\mid{\bm{Q}}\Bigr{)}.$
@@
Let us remember that because of $Q_{1}^{*}+Q_{2}^{*}+Q_{3}^{*}=1$, we have
@@ A.1 The explicit expressions of $Q_{h}^{*}$ and $\chi_{h}$ at $r\to 0$
$Q_{2}^{*}(T,0)=\frac{\exp(-2\beta)}{2+\exp(-2\beta)}.$ (40)
@@
There are six cases in terms of the possibilities of taking values of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$, each of which can be $-1$, 00, or $1$. Note that if one of $G_{\beta}^{(1)}(a)$ and $G_{\beta}^{(3)}(a)$ is 00, the value of the other is uniquely determined. Indeed, when the two conditions of
$-2Q_{1}^{*}+2Q_{3}^{*}+1>0,$ (42)
$2Q_{1}^{*}-2Q_{3}^{*}+1>0,$ (43)
@@
for $h=1,3$. Additionally, the two conditions of (42) and (43) lead to
@@
$Q_{3}^{*}(0,0)=1-Q_{1}^{*}(0,0),$ (47)
$Q_{2}^{*}(0,0)=0.$ (48)
Note that the other five cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of Eq. (36).
@@
Then, substituting this form with (39) and (40) into (36), for $O(\beta r)\ll 1$ up to $O(r)$ order, we obtain
@@
$\chi_{3}(T,0)=4\frac{-4g_{\beta}(2+g_{\beta})+4g_{\beta}^{\prime}\beta(2-g_{\beta})}{(2+g_{\beta})^{2}((2+g_{\beta})(2-g_{\beta})+8g_{\beta}^{\prime}\beta)},$ (54)
@@
A.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
Among six possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, let us consider the situation where the following two conditions hold for $T\ll 1$:
$-2Q_{1}^{*}+2Q_{3}^{*}+1=0,$ (55)
$0\displaystyle 2Q_{1}^{*}-2Q_{3}^{*}+1>$ 0 . (56)
@@ A.2 The explicit expressions of $Q_{h}^{*}$ and $\mu_{h}$ at $T\to 0$
Combining (59) with the condition of (55) gives
@@
$Z_{3}=$ $Q_{3}^{*}(0,r)=-\frac{1}{4}r+\frac{1}{4},$ (62)
$Q_{2}^{*}(0,r)=\frac{r}{2}.$ (63)
Let us move onto (58) into which we substitute (61) and (62), leading to
@@
Since $-1<\tanh 2(-\mu_{1}+\mu_{3})<1$ should hold, $r$ satisfies $0<r<r_{0}=-3+2\sqrt{3}$ due to $\lim_{r\to 0}\theta(r)=1$ and $\lim_{r\to r_{0}}\theta(r)=-1$. Thus, it turns out that (61), (62) and (63) as the stationary solution is valid only for $0<r<r_{0}$. Note that $\lim_{r\to 0}Q_{1}^{*}(0,r)=\sup Q_{1}^{*}(0,0)=\dfrac{3}{4}$, $\lim_{r\to 0}Q_{3}^{*}(0,r)=\inf Q_{3}^{*}(0,0)=\dfrac{1}{4}$, and $\lim_{r\to 0}Q_{2}^{*}(0,r)=Q_{2}^{*}(0,0)=0$ hold.
Next, among the other five possibilities of taking values of $G_{\beta}^{(1)}$ and $G_{\beta}^{(3)}$, when the following two conditions
$-2Q_{1}^{*}+2Q_{3}^{*}+1<0,$ (65)
$2Q_{1}^{*}-2Q_{3}^{*}+1>0,$ (66)
@@
$Q_{3}^{*}(0,r)=\frac{(1-r)^{2}}{(1+r)^{2}},$ (72)
$Q_{2}^{*}(0,r)=\frac{2r(1-r)}{(1+r)^{2}}.$ (73)
Substituting (71) and (72) into the two conditions of (65) and (66) gives
$3r^{2}+10r-1>0,$ (74)
$r^{2}+6r-3>0,$ (75)
which lead to $r_{0}<r\leq 1$. Thus, the obtained solution of (71), (72), and (73) is valid only for $r_{0}<r\leq 1$. Combining two cases of smaller $r$ and larger $r$ than $r_{0}$, the left-sided limit and the right-sided limit of $Q_{h}^{*}$ as $r\to r_{0}$ coincide. Note that the other four cases for the possibilities of taking values of $G_{\beta}^{(h)}(a)$ give no solutions of Eq. (36).
@@
$\mu_{3}(0,r)=\frac{1-r}{4}\text{artanh }\theta(r).$ (80)
@@
A.3 The explicit expressions of $Q_{h}^{*}$ and $\nu_{h}$ at $\beta\to 0$
@@
$Q_{3}^{*}(\infty,r)=\frac{(1-r)^{2}}{3r^{2}+10r+3},$ (88)
$Q_{2}^{*}(\infty,r)=\frac{-3r^{2}+2r+1}{3r^{2}+10r+3}.$ (89)
@@
Note that by taking the derivative of the stationary solutions from (87), (88), and (89) with respect to $r$, we also have
@@
$\chi_{2}(\infty,0)=-\frac{4}{9},$ (96)
which are consistent with the expression of (53) and (54) in the limit of $\beta\to 0$.
Appendix B Monte Carlo sampling
Figure 8: The density of height, $\sum_{i}\delta_{x_{i},h}/N$, calculated by the Monte Carlo (MC) simulation with $N=1000$ and the solutions $Q_{h}$ of the population dynamics (33). $T=1$, $r=0.2$.
B.1 Updating rules
Figure 9: The variance $\sigma_{hh}$ and correlation coefficients $C_{hh^{\prime}}$ averaged over ten samples with different trajectories. $T=10^{-3}$, $t_{0}=100$, $\tau=150$, $N=10^{2},10^{3},10^{4}$.
@@ B.1 Updating rules
  2. (1)
    Randomly choose a name $\ell\ (1\leq\ell\leq N)$ of an agent with uniform distribution. That is, the probability of taking each name is $1/N$.
  3. (2)
@@
  4. (3)
    If the chosen agent $\ell$ is not disturbed, choose either a positive or negative direction with probability $1/2$.
@@
      If a positive direction is chosen and $x_{\ell}\neq 3$, make the transition from state $x_{\ell}$ to $x_{\ell}+1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh{\beta\Big{(}2(x_{\ell}-\frac{1}{N-1}\sum_{k\neq\ell}x_{k})}+1\Big{)}\Bigg{)}.$ (97)
@@
      If a negative direction is chosen and $x_{\ell}\neq 1$, make the transition from state $x_{\ell}$ to $x_{\ell}-1$ with the probability
      $\frac{1}{2}\Bigg{(}1+\tanh\beta\Big{(}-2(x_{\ell}-\frac{1}{N-1}\sum_{k\neq\ell}x_{k})+1\Big{)}\Bigg{)}.$ (98)
As shown in Fig. 8, we have confirmed that Monte Carlo simulations have good agreement with the solution of the population dynamics (33). Note that for convenience in the simulations, we have replaced $\frac{1}{N-1}\sum_{k\neq\ell}x_{k}$ by $\frac{1}{N}\sum_{k}x_{k}$; the caused errors are the order of $1/N$, which is negligible, at least, near stationary states for sufficiently large $N$, except for the vicinity of $r=1,T=0$.
B.2 Finite-size fluctuations
@@
As shown in Fig. 9, we have performed Monte Carlo simulations with different systems sizes $N$. If we straightforwardly extrapolate the behaviors for larger system sizes, there seem to be singular behaviors in the three types of variance and the three types of correlation function in the large size limit near $r=r_{0}$ for $r<r_{0}$. In particular, it is notable to point out that three correlation coefficients in the large size limit seem to have perfect correlation, corresponding to $|C_{hh^{\prime}}|=1$, for only $r<r_{0}$.
Additionally, the results obtained by Monte Carlo simulations suggest that $C_{hh^{\prime}}$ for $r<r_{0}$ tends to show U-shaped curves as the system size $N$ gets larger. It is notable that the bottoms of the U-shaped curves are close to $r=-5+2\sqrt{7}\simeq 0.291\cdots$, which is the point where $\mu_{h}$ changes its sign. Nevertheless, more works with sufficiently large samples need to be done to obtain more convincing data, which remain to be found as another work in the future.